# UniVS: Unified and Universal Video Segmentation with Prompts as Queries

### v1

@@ 5.1 Experimental Settings
Datasets. The VIS datasets include YouTube-VIS 2019/2021 (YT19/21) [yang2019video] and OVIS [qi2021occluded]. The VSS and VPS datasets respectively include VSPW [vspw] and VIPSeg. The VOS datasets include DAVIS [davis], YouTube-VOS 2018 (YT18) [youtubevos], MOSE [mose] and BURST [burst]. RefVOS datasets include RefDAVIS [refvos] and Ref-YouTube-VOS (RefYT) [refvos], while PVOS uses VIPOSeg [xu2023viposeg] as dataset. The 2017 version of DAVIS and RefDAVIS are used [davis]. Note that VSPW, VIPSeg and VIPOseg share the same original videos but have different protocols and guidelines for annotation. Due to the high cost of annotating datasets, the video scenes contained in each dataset lack diversity, being mostly sports scenes or animal scenes.
Implementation Details. We use Detectron2 [wu2019detectron2] and follow Mask2Former [cheng2021mask2former] baseline settings. For data augmentation, we use the large-scale jittering (LSJ) augmentation [du2021lsj_aug] with a random scale sampled from range 0.5 to 4.0, followed by fixed-size cropping and padding to $1024\times 1024$. We use the distributed training framework with 16 V100/A100 GPUs. Each mini-batch has 3 images or 1 video clip (2-7 frames) per GPU. ResNet50 [he2016resnet] and SwinT/B/L [swin] are adopted as the backbone networks, while we use the CLIP Text Encoder, whose corresponding visual encoder is ResNet50x4. We sample 32 points per visual prompt for visual prompt-guided targets. Considering the difference in the number of objects on various videos, we increase the number of learnable queries from 100 to 200 in the unified video mask decoder.
@@
  Video Tasks VIS VSS VPS
Method
Back-
YT19
YT21
OVIS
VSPW VIPSeg
bone
mAP
mAP
mAP
mIoU

### v2

@@ 5.1 Experimental Settings
Datasets. The VIS datasets include YouTube-VIS 2019/2021 (YT19/21) [87] and OVIS [63]. The VSS and VPS datasets respectively include VSPW [57] and VIPSeg. The VOS datasets include DAVIS [62], YouTube-VOS 2018 (YT18) [84], MOSE [16] and BURST [2]. RefVOS datasets include RefDAVIS [68] and Ref-YouTube-VOS (RefYT) [68], while PVOS uses VIPOSeg [85] as dataset. The 2017 version of DAVIS and RefDAVIS are used [62]. Note that VSPW, VIPSeg and VIPOseg share the same original videos but have different protocols and guidelines for annotation. Due to the high cost of annotating datasets, the video scenes contained in each dataset lack diversity, being mostly sports scenes or animal scenes.
Implementation Details. We use Detectron2 [82] and follow Mask2Former [9] baseline settings. For data augmentation, we use the large-scale jittering (LSJ) augmentation [17] with a random scale sampled from range 0.5 to 4.0, followed by fixed-size cropping and padding to $1024\times 1024$. We use the distributed training framework with 16 V100/A100 GPUs. Each mini-batch has 3 images or 1 video clip (2-7 frames) per GPU. ResNet50 [23] and SwinT/B/L [51] are adopted as the backbone networks, while we use the CLIP Text Encoder, whose corresponding visual encoder is ResNet50x4. We sample 32 points per visual prompt for visual prompt-guided targets. Considering the difference in the number of objects on various videos, we increase the number of learnable queries from 100 to 200 in the unified video mask decoder.
@@
     Video Tasks VIS VSS VPS
Method Back- YT19 YT21 OVIS VSPW VIPSeg
bone mAP mAP mAP mIoU mVC8 VPQ STQ
IDOL[80] R50 49.5 43.9 30.2 - - - -
MinVIS[26] R50 47.4 44.2 25.0 - - - -
MDQE[38] R50 * 44.5 29.2 - - - -
GenVIS-off[24] R50 51.3 46.3 34.5 - - - -
TCB[57] R101 - - - 37.8 87.8 - -
ClipPanoFCN[56] R50 - - - - - 22.9 31.5
IDOL[80] SwinL 49.5 43.9 42.6 - - - -
MinVIS[26] SwinL - 55.3 41.6 - - - -
GenVIS-off[24] SwinL 63.8 60.1 45.4 - - - -
CFFM[70] MiTB5 - - - 49.3 90.8 - -
@@
     Video Tasks VOS RefVOS PVOS
Method Back- DAVIS YT18 RefDAVIS RefYT VIPOSeg
bone $G^{th}$ $G^{th}$ J&F J&F $G^{th\&sf}$
XMem[12] R50 86.2 85.7 - - -
DeAOT[89] R50 85.2 86.0 - - -
ReferFormer[78] R50 - - 58.5 55.6 -
OnlineRefer[76] R50 - - 59.3 57.3 -
SgMg[55] SwinT - - 61.9 62.0 -
PAOT[85] R50 - - - - 75.4
DeAOT[89] Swin-B 86.2 86.2 - - -
OnlineRefer[76] Swin-L - - 64.8 63.0 -
SgMg[55] SwinB - - 63.3 65.7 -
PAOT[85] SwinB - - - - 75.3
(b) Prompt-specified VS models
Table 1: Quantitative performance comparison on individual VS models, which are specifically designed based on the characteristics of each task. ‘-’ means the model lacks this capability. The best results are shown in bold.
       Video Tasks VIS VSS VPS VOS RefVOS PVOS
Method Backbone Joint Universal YT19 YT21 OVIS VSPW VIPSeg DAVIS YT18 DAVIS RefYT VIPOSeg
Training mAP mAP mAP mIoU mVC8 VPQ STQ $G^{th}$ $G^{th}$ J&F J&F $G^{th\&sf}$
VideoK-Net[44] ResNet50 40.5 * * * * 26.1 33.1 - - - - -
Video-kMax[69] ResNet50 * * * 44.3 86.0 38.2 39.9 - - - - -
Tube-Link[43] ResNet50 52.8 47.9 29.5 42.3 86.8 39.2 39.5 - - - - -
DVIS-off[95] ResNet50 52.6 47.4 33.8 * * 43.2 42.8 - - - - -
UniRef[79] ResNet50 $\checkmark$ - - - - - - - * 81.4 63.5 60.6 *
TarVIS[1] ResNet50 $\checkmark$ * 48.3 31.1 * * 33.5 43.1 82.6 * - - -
UNINEXT[86] ResNet50 $\checkmark$ 53.0 * 34.0 - - - - 74.5 77.0 63.9 61.2 -
UniVS ResNet50 $\checkmark$ $\checkmark$ 47.4 46.6 30.8 48.2 88.5 38.6 45.8 70.5 69.2 57.9 56.2 60.2
UniVS SwinT $\checkmark$ $\checkmark$ 52.4 51.6 33.0 51.3 89.4 38.9 51.7 71.7 70.3 58.5 56.2 62.3
TubeFormer[30] A-R50$\times$64 47.5 41.2 * 63.2 92.1 * * - - - - -
VideoK-Net+[44] SwinB 51.4 * * * * 39.8 46.3 - - - - -
Video-kMax[69] ConvNeXtL * * * 63.6 91.8 51.9 51.7 - - - - -
Tube-Link[43] SwinL 64.6 58.4 * 59.7 90.3 * * - - - - -
DVIS-off[95] SwinL 64.9 60.1 49.9 * * 57.6 55.3 - - - - -
UniRef[79] SwinL $\checkmark$ - - - - - - - * 82.6 66.3 67.4 *
TarVIS[1] SwinL $\checkmark$ * 60.2 43.2 * * 48.0 52.9 85.2 * - - -
UNINEXT[86] ConvNeXtL $\checkmark$ 64.3 * 41.1 - - - - 77.2 78.1 66.7 66.2 -
UNINEXT[86] ViT-H $\checkmark$ 66.9 * 49.0 - - - - 81.8 78.6 72.5 70.1 -
UniVS SwinB $\checkmark$ $\checkmark$ 57.8 56.5 39.0 59.4 90.4 46.7 56.1 75.0 70.9 58.6 57.4 68.2
UniVS SwinL $\checkmark$ $\checkmark$ 60.0 57.9 41.7 59.8 92.3 49.3 58.2 76.2 71.5 59.4 58.0 68.6
Table 2: Overall quantitative performance comparison of unified VS models. ‘-’ means the model lacks this capability and ‘*’ means the result is not reported. The results of UniVS with SwinT backbone and UNINEXT with ViT-H backbone are also listed in gray color. However, due to use of different backbones, they are not considered in the performance comparison. The best results are shown in bold.
5.2 Video Benchmark Results
We compare the results of recent high-performance VS models. The results of individual models trained on a single task/dataset are shown in Table 1, and the results of unified models trained individually or jointly on different tasks are shown in Table 2. To assess the generalization capability of competing models, we present the quantitative performance comparison on 10 benchmarks of six VS tasks, including VIS, VSS, VPS, VOS, RefVOS and PVOS. The detailed quantitative comparison on more image/video benchmarks can be found in the Supplementary Material.
Individual models are specifically designed based on the characteristics of each task, resulting in high performance on each benchmark, as shown in Table 1. For VIS task, GenVIS-off [24] with sequential learning achieves impressive performance of 51.3, 46.3 and 34.5 mAP on YT19, YT21 and OVIS, respectively. For the recently proposed VSS/VPS tasks, the baselines [57, 29] only obtain decent performance. For VOS task, XMem[12] and DEAOT [89] employ long-term information propagation modules, achieving state-of-the-art performance of 86.2 on DAVIS and 86.0 on YT18, respectively. For RefVOS task, SgMg [55] achieves the best performance of 62.0 on RefYT by using a spectrum-guided multi-granularity approach. For the newly proposed PVOS task, PAOT [85] extends the typical VOS method DEAOT, obtaining 75.4 on VIPOSeg.
Unified models trained individually on different tasks include VideoK-Net [44], Video-kMax [69], Tube-Link [43], DVIS [95] and TubeFormer [30], as listed in Table 2, which tend to handle both thing and stuff categories within a single framework. Video-kMax with clip k-means mask transformer achieves high performance on VSS task, reaching 44.3 mIoU on VSPW. By using the strong Mask2Former[9] as baseline, Tube-Link and DVIS inherit its ability to handle both thing and stuff classes, achieving remarkable performance on VIS and VPS tasks, as evidenced by the 33.8 mAP on OVIS and 43.2 VPQ on VIPSeg. It should be noted that the above unified models are trained individually on each dataset so that they can achieve excellent performance on each benchmark; however, this makes them lack the generalization ability to other datasets. Additionally, they cannot handle prompt-specified VS tasks, including VOS, RefVOS and PVOS.
Unified models trained jointly on different tasks aim to accommodate as many VS tasks as possible within a single model, where all tasks can be accomplished using the same set of trained weights. As shown in the top rows of Table 2, UniRef [79] unifies all prompt-specified VS tasks, but it fails to deal with category-specified VS tasks, such as VIS, VSS and VPS. TarVIS [1] can handle most VS tasks except for RefVOS due to the lack of language encoding capabilities. UNINEXT [86] focuses solely on object-centric segmentation without considering entity segmentation of stuff categories, thereby being incapable of handling VSS, VPS and PVOS tasks. In contrast, our UniVS is the only approach that accommodates all VS tasks within a single model, demonstrating the highest generalization capability in universal segmentation.
Specifically, UniRef achieves outstanding performance on VOS and RefVOS tasks, reaching 81.4 on YT18. TarVIS achieves state-of-the-art performance on datasets with simple scenes, such as 48.3 mAP on YT21 of VIS task and 82.6 on DAVIS of VOS task. UNINEXT utilizes a large amount of image/video data to jointly train the model, allowing it to achieve state-of-the-art performance in instance-level segmentation benchmarks, such as 53.0/34.0 mAP on YT19/OVIS for VIS task and 63.9/61.2 on DAVIS and RefYT for RefVOS task. Our UniVS achieves comparable performance on VIS task, top-ranked performance on VSS/VPS tasks and slightly lower performance in prompt-specified VS tasks. For VSS and VPS tasks, UniVS brings 2$\sim$4performance improvement, reaching 48.2 mIoU and 88.5 mVC8 on VSPW and 45.8 STQ on VIPSeg.
Compared with UniRef, TarVIS and UNINEXT, our UniVS can handle category-specified VS tasks, text prompt-guided segmentation and entity segmentation for stuff categories simultaneously. Pursuing a more universal segmentation capability, however, may sacrifice performance in individual tasks due to task conflicts and limited video data. Compared to UniRef, which utilizes heavy temporal propagation modules to pass reference frames and masks to the current frame, UniVS adopts a parameter-free Visual Sampler to extract prompt information, resulting in a slight decrease in performance. Compared to UNINEXT, which uses 900 object queries, we train UniVS using only 200 object queries with fewer video data. Additionally, UniVS needs to accommodate stuff categories in object segmentation. These two factors explain the slightly lower performance of our approach in instance-level segmentation benchmarks compared to UNINEXT. In future work, more diverse training video data and long-term information propagation modules will be explored to improve the performance of our unified segmentation models.
Our results with stronger backbones further validate the aforementioned conclusion, as shown in the bottom rows of Tables 1 and 2. Our UniVS is the only model that is capable of handling all the six VS tasks within a single framework, setting new state-of-the-art performance on VSS and VPS tasks, and achieving 92.3 mVC8 on VSPW and 58.2 STQ on VIPSeg. Overall, our UniVS can obtain an appropriate balance between performance and universality capability.
     Video Tasks VIS VPS VOS RefVOS